# Information Inequalities for Five Random Variables

### v1

@@ 4. What to compute? How to compute?
As discussed in Section 3, the task of finding new non-Shannon entropy inequalities implied by the existence of an $n,m$-copy reduces to enumerating all facets of the polyhedral cone $\mathcal{Q}^{*}$ defined in (2). However, without further reduction, this polyhedral computation is intractable even for small parameter values. Therefore, in this Section we look at some general methods to reduce the complexity of the computation, and then discuss how the number of elements in the $Y\mkern-2.0muXZ$ partition were chosen.
@@ 4.1. Tight and modular parts
When $\alpha$ is set to $f(i\mkern 1.0mu{|}\mkern 1.0muN\mkern 1.5mu{-}\mkern 1.5mui)$, $f\mathord{\downarrow}^{i}_{\alpha}$ is denoted simply by $f\mathord{\downarrow}^{i}$. Note that for $i\notin A$ we have $f(Ai)-f(A)\geqslant f(N)-f(N\mkern 1.5mu{-}\mkern 1.5mui)=f(i\mkern 1.0mu{|}\mkern 1.0muN\mkern 1.5mu{-}\mkern 1.5mui)$ by submodularity. Consequently, $f\mathord{\downarrow}^{i}$ can be written explicitly as
@@
Therefore $f=f\mathord{\downarrow}^{i}+f(i\mkern 1.0mu{|}\mkern 1.0muN\mkern 1.5mu{-}\mkern 1.5mui)\mkern 1.5mu\mathbfit r_{i}$, where $\mathbfit r_{i}$ is the polymatroid defined in (3). The result of tightening $f$ at $i$ is the function $f\mathord{\downarrow}^{i}$. The tight part of $f$, denoted by $f\mathord{\downarrow}$, is the result of tightening $f$ at every element of its base set $N=\{i_{1},\dots,i_{n}\}$:
@@
$f=f\mathord{\downarrow}+\sum_{i\in N}f(i\mkern 1.0mu{|}\mkern 1.0muN\mkern 1.5mu{-}\mkern 1.5mui)\mkern 1.5mu\mathbfit r_{i}.$
@@
From the above it follows that the facets of the cone $\mathcal{Q}^{*}$ belong to two disjoint groups. There are $|N|$ (trivial, Shannon) facets that bound the modular part of $\mathcal{Q}^{*}$, and the rest that bound the tight part. The normal vectors of the facets in the second group are balanced, and only they can provide non-Shannon inequalities. Therefore, it suffices to consider only the tight part of $\mathcal{Q}^{*}$. This part is generated by a smaller collection of polymatroids, has fewer dimensions, and so can be handled more efficiently.
@@
The tight part of $\mathcal{Q}^{*}$ is generated by the $n,m$-copies of the polymatroids $f$ on $N=Y\mkern-2.0muXZ$ that are (i) tight; (ii) satisfy $f(Y,Z\mkern 1.0mu{|}\mkern 1.0muX)=0$; and (iii) for all $y\in Y$, $f(y\mkern 1.0mu{|}\mkern 1.0muY\mkern-2.0muX\mkern 1.5mu{-}\mkern 1.5muy)=0$, and for all $z\in Z$, $f(z\mkern 1.0mu{|}\mkern 1.0muXZ\mkern 1.5mu{-}\mkern 1.5muz)=0$.
Observe that the tightness of $f$ at the elements of $Y$ and $Z$ follows from condition (iii) and submodularity, thus (i) is relevant only for elements of $X$.
@@
Let $f^{*}$ be an $n,m$-copy of $f$. In the definition of $\mathcal{Q}^{*}$ only the values of $f\mathord{\upharpoonright}Y\mkern-2.0muX$ and the values of $f\mathord{\upharpoonright}XZ$ are used. Therefore, $f$ can be replaced with any other polymatroid that has the same restrictions. Such a polymatroid is $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ by part (i) of Definition 1, which gives (ii). For (iii) let $y\in Y$, and $\alpha=f(y\mkern 1.0mu{|}\mkern 1.0muY\mkern-2.0muX\mkern 1.5mu{-}\mkern 1.5muy)$. Apply Lemma 3 to $f^{*}$ and all instances of $y$ in the copies $Y_{i}$ to get the new polymatroid $g^{*}$. Denoting the instance of $y$ in $Y_{1}$ by $y_{1}$, the lemma provides $g^{*}(y_{1}\mkern 1.0mu{|}\mkern 1.0muY_{1}X\mkern 1.5mu{-}\mkern 1.5muy_{1})=0$. In addition, $g^{*}$ is an $n,m$-copy of its restriction to $Y_{1}XZ_{1}$. Since this restriction and the polymatroid $f$ differ only by a modular shift on subsets of $Y\mkern-2.0muX$ and $XZ$, their tight parts are the same. A similar reduction on elements of $Z$, and finally on elements of $X$ provide the statement. ∎
@@ 4.3. No new inequality
Suppose that the polymatroids $(f_{1},Y\mkern-2.0muX)$ and $(f_{2},XZ)$ have a common restriction on $X$ which is modular. Then there is a polymatroid $(g,Y\mkern-2.0muXZ)$ that extends both $f_{1}$ and $f_{2}$ such that $g(Y,Z\mkern 1.0mu{|}\mkern 1.0muX)=0$.
@@
Using the fact that $f_{1}\mathord{\upharpoonright}X$ and $f_{2}\mathord{\upharpoonright}X$ are isomorphic and modular, a simple calculation shows that $g$ is a polymatroid, and satisfies the requirements. For details, consult [10, 12] or [26]. ∎
@@
We claim that the cone generated by the tight part of $n,m$-copies is the same as the cone generated by the $n,1$-copies. To prove this, let $f$ be a polymatroid on $Y\mkern-2.0muXZ$ that satisfies the conditions of Claim 4, and let $f^{*}$ be an $n,1$-copy of $f$ so that $f$ is identified with $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ where $Z_{1}$ has a single element $z$. Let $g^{*}$ be the polymatroid when $m\mkern 1.5mu{-}\mkern 1.5mu1$ identical copies of $z$ are added to $f^{*}$. We claim that $g^{*}$ is an $n,m$-copy. The only non-trivially satisfied condition is that the copies of $z$, $z_{1}$ and $z_{2}$, are independent over $X$. Since $f(z\mkern 1.0mu{|}\mkern 1.0muX)=0$ by (iii) of Claim 4, we have $g^{*}(Xz_{1})=g^{*}(Xz_{2})=g^{*}(Xz_{1}z_{2})=g^{*}(X)$, thus $g^{*}(z_{1},z_{2}\mkern 1.0mu{|}\mkern 1.0muX)=0$. Since $g^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ and $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ are the same polymatroids, the $n,m$-cone is part of the $n,1$-cone, as claimed. ∎
@@ 4.4. Problem Parameters
By Claim 6, the Maximum Entropy Method does not yield new inequalities when $f\mathord{\upharpoonright}X$ is modular. This is certainly the case when $|X|=1$, so we must have $|X|\geqslant 2$. By Claim 8, if $|Y|=|Z|=1$, then beyond the $1,1$-copy, no additional inequalities are generated. The smallest parameter setting when new entropy inequalities are be expected as the number of copies grows is $|Y|=2$, $|X|=2$ and $|Z|=1$. We fix these sizes, as well as the labels of the members of each set as
@@
    Fix the number of copies $n$, called a generation. Determine the generating matrix $M$ of the cone $\mathcal{Q}^{*}$ as specified in Claim 2 using only polymatroids that satisfy the conditions of Claim 4.
@@ 5. Computation
The cone $\mathcal{Q}^{*}$ whose non-Shannon bounding facets provide the new entropy inequalities sits in the $d_{1}=19$-dimensional Euclidean space with coordinates indexed by the non-empty subsets of $Y\mkern-2.0muX=\{abcd\}$ and $XZ=\{abz\}$. Fix the number of copies to $n\geqslant 1$. This choice also fixes the dimension $d_{3}$ of the vector $\mathbf{y}$. The generating matrix $M$ of the polyhedral cone $\mathcal{Q}^{*}$ from (2) is repeated here:
$\mathcal{Q}^{*}=\big\{\mathbf{x}\in\mathbb{R}^{19}:M\mkern 1.5mu{\cdot}\mkern 1.5mu(\mathbf{x},\mathbf{y})^{\top}\geqslant 0\mbox{ for some }\mathbf{y}\in\mathbb{R}^{d_{3}}\big\}.$ (4)
The modular part of $\mathcal{Q}^{*}$ is $5$-dimensional, and so its tight part sits in a $14$-dimensional subspace of $\mathbb{R}^{19}$.
A structural property of the polymatroid region on the four-element set abcdabcd allows us to further reduce the complexity of the polyhedral computation required in step (2) above. This region has a central part and six permutationally equivalent “protrusions,” depending on the signs of the Ingleton expressions
@@
If all of them are non-negative, then the restriction $f\mathord{\upharpoonright}abcd$ is a linear polymatroid; otherwise exactly one of these Ingleton expressions is negative, see e.g. [25]. Accordingly, the cone $\mathcal{Q}^{*}$ is cut into seven parts by these Ingleton hyperplanes: the central part where all Ingleton values are non-negative, and six other parts where exactly one of the expressions is negative. The facets of each part can be computed separately.
Parts of $\mathcal{Q}^{*}$ on the negative side of $\mkern 1.5mu], [adbc][adbc$], $bcad\mkern 1.5mu]$, and $[bdac]$ are isomorphic because swapping $a\leftrightarrow b$ and/or $c\leftrightarrow d$ are symmetries of $\mathcal{Q}^{*}$. Therefore, it suffices to consider only one of them. The central part, where every Ingleton expression is non-negative, does not yield new inequalities. This follows from Lemma 9 below, as the elements of the central part are linear.
@@
Since every polymatroid on $3$ elements is linear, and linearly representable polymatroids on three or four elements are representable over any field, we can assume, after scaling and using continuity, that both $f\mathord{\upharpoonright}abcd$ and $f\mathord{\upharpoonright}abz$ are $\mathbb{F}$-linearly representable over the same finite field $\mathbb{F}$. Denote the two representing vector spaces by $V^{1}$ and $V^{2}$, and consider the subspace arrangements $(V^{1}_{a},V^{1}_{b})$ and $(V^{2}_{a},V^{2}_{b})$ in the two vector spaces. Now $V^{i}_{a}$ and $V^{i}_{b}$ have dimensions $f(a)$ and $f(b)$, respectively, and their linear span has dimension $f(ab)$. Therefore, these arrangements are isomorphic, and $V^{1}$ and $V^{2}$ can be glued along the linear span of $(V^{1}_{a},V^{1}_{b})$ and $(V^{2}_{a},V^{2}_{b})$. This gluing yields an $\mathbb{F}$-linear polymatroid $g$ that has the same restrictions on abcdabcd and on abzabz as $f$ does. Since this $g$ is entropic, it has an $n$-copy for every $n\geqslant 1$. This $n$-copy is also an $n$-copy of $f$, as required. ∎
Consequently, up to the $a\leftrightarrow b$ and $c\leftrightarrow d$ symmetries, three mutually exclusive cases are left: $f[abcd\mkern 1.5mu]<0$, $f[acbd\mkern 1.5mu]<0$, and $f[cdab\mkern 1.5mu]<0$. Using the homogeneity of $\mathcal{Q}^{*}$, the Ingleton value can be set to 1-1, in effect taking a cross-section of $\mathcal{Q}^{*}$ that has one fewer dimensions. Facets of the part of $\mathcal{Q}^{*}$ we are considering are also facets of these cross-sections; consequently, only facets of the cross-sections need to be computed. We consider these three cases separately in the subsections below.
@@
  • apart from the Ingleton coordinate, other coordinates have non-negative values.
The first property allows to set the Ingleton value explicitly. Based on the second property, the tight part of the cross-section can be separated by dropping some coordinates; and the third property potentially reduces the complexity of the polyhedral enumeration algorithm.

### v2

@@ 4. What to Compute? How to Compute?
As discussed in Section 3, the task of finding new non-Shannon entropy inequalities implied by the existence of an $n,m$-copy reduces to enumerating all facets of the polyhedral cone $\mathcal{Q}^{*}$ defined in (2). However, without further reduction, this polyhedral computation is intractable even for small parameter values. Therefore, in this Section we look at some general methods to reduce the complexity of the computation, and then discuss how the number of elements in the $Y\mkern-2.0muXZ$ partition was chosen.
@@ 4.1. Tight and Modular Parts
When $\alpha$ is set to $f(i\mkern 1.5mu{|}\mkern 1.5muN\mkern 1.5mu{-}\mkern 1.5mui)$, $f\mathord{\downarrow}^{i}_{\alpha}$ is denoted simply by $f\mathord{\downarrow}^{i}$. Note that for $i\notin A$ we have $f(Ai)-f(A)\geqslant f(N)-f(N\mkern 1.5mu{-}\mkern 1.5mui)=f(i\mkern 1.5mu{|}\mkern 1.5muN\mkern 1.5mu{-}\mkern 1.5mui)$ by submodularity. Consequently, $f\mathord{\downarrow}^{i}$ can be written explicitly as
@@
Therefore, $f=f\mathord{\downarrow}^{i}+f(i\mkern 1.5mu{|}\mkern 1.5muN\mkern 1.5mu{-}\mkern 1.5mui)\mkern 1.5mu\mathbfit r_{i}$, where $\mathbfit r_{i}$ is the polymatroid defined in (3). The result of tightening $f$ at $i$ is the function $f\mathord{\downarrow}^{i}$. The tight part of $f$, denoted by $f\mathord{\downarrow}$, is the result of tightening $f$ at every element of its base set $N=\{i_{1},\dots,i_{n}\}$:
@@
$f=f\mathord{\downarrow}+\sum_{i\in N}f(i\mkern 1.5mu{|}\mkern 1.5muN\mkern 1.5mu{-}\mkern 1.5mui)\mkern 1.5mu\mathbfit r_{i}.$
@@
From the above it follows that the facets of the cone $\mathcal{Q}^{*}$ belong to two disjoint groups. There are $|N|$ (trivial, Shannon) facets that bound the modular part of $\mathcal{Q}^{*}$, and the rest bound the tight part. The normal vectors of the facets in the second group are balanced, and only they can provide non-Shannon inequalities. Therefore, it suffices to consider only the tight part of $\mathcal{Q}^{*}$. This part is generated by a smaller collection of polymatroids, has fewer dimensions, and so can be handled more efficiently.
@@
The tight part of $\mathcal{Q}^{*}$ is generated by the $n,m$-copies of the polymatroids $f$ on $N=Y\mkern-2.0muXZ$ that are (i) tight; (ii) satisfy $f(Y,Z\mkern 1.5mu{|}\mkern 1.5muX)=0$; and (iii) for all $y\in Y$, $f(y\mkern 1.5mu{|}\mkern 1.5muY\mkern-2.0muX\mkern 1.5mu{-}\mkern 1.5muy)=0$, and for all $z\in Z$, $f(z\mkern 1.5mu{|}\mkern 1.5muXZ\mkern 1.5mu{-}\mkern 1.5muz)=0$.
Observe that the tightness of $f$ at the elements of $Y$ and $Z$ follows from condition (iii) and submodularity; thus (i) is relevant only for elements of $X$.
@@
Let $f^{*}$ be an $n,m$-copy of $f$. In the definition of $\mathcal{Q}^{*}$ only the values of $f\mathord{\upharpoonright}Y\mkern-2.0muX$ and the values of $f\mathord{\upharpoonright}XZ$ are used. Therefore, $f$ can be replaced with any other polymatroid that has the same restrictions. Such a polymatroid is $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ by part (i) of Definition 1, which gives (ii). For (iii) let $y\in Y$, and $\alpha=f(y\mkern 1.5mu{|}\mkern 1.5muY\mkern-2.0muX\mkern 1.5mu{-}\mkern 1.5muy)$. Apply Lemma 3 to $f^{*}$ and all instances of $y$ in the copies $Y_{i}$ to get the new polymatroid $g^{*}$. Denoting the instance of $y$ in $Y_{1}$ by $y_{1}$, the lemma provides $g^{*}(y_{1}\mkern 1.5mu{|}\mkern 1.5muY_{1}X\mkern 1.5mu{-}\mkern 1.5muy_{1})=0$. In addition, $g^{*}$ is an $n,m$-copy of its restriction to $Y_{1}XZ_{1}$. Since this restriction and the polymatroid $f$ differ only by a modular shift on subsets of $Y\mkern-2.0muX$ and $XZ$, their tight parts are the same. A similar reduction on elements of $Z$, and finally on elements of $X$ provide the statement. ∎
@@ 4.3. No New Inequality
Suppose that the polymatroids $(f_{1},Y\mkern-2.0muX)$ and $(f_{2},XZ)$ have a common restriction on $X$ which is modular. Then there is a polymatroid $(g,Y\mkern-2.0muXZ)$ that extends both $f_{1}$ and $f_{2}$ such that $g(Y,Z\mkern 1.5mu{|}\mkern 1.5muX)=0$.
@@
Using the fact that $f_{1}\mathord{\upharpoonright}X$ and $f_{2}\mathord{\upharpoonright}X$ are isomorphic and modular, a simple calculation shows that $g$ is a polymatroid and satisfies the requirements. For details, consult [10, 12] or [26]. ∎
@@
We claim that the cone generated by the tight part of $n,m$-copies is the same as the cone generated by the $n,1$-copies. To prove this, let $f$ be a polymatroid on $Y\mkern-2.0muXZ$ that satisfies the conditions of Claim 4, and let $f^{*}$ be an $n,1$-copy of $f$ so that $f$ is identified with $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ where $Z_{1}$ has a single element $z$. Let $g^{*}$ be the polymatroid when $m\mkern 1.5mu{-}\mkern 1.5mu1$ identical copies of $z$ are added to $f^{*}$. We claim that $g^{*}$ is an $n,m$-copy. The only non-trivially satisfied condition is that the copies of $z$, $z_{1}$ and $z_{2}$, are independent over $X$. Since $f(z\mkern 1.5mu{|}\mkern 1.5muX)=0$ by (iii) of Claim 4, we have $g^{*}(Xz_{1})=g^{*}(Xz_{2})=g^{*}(Xz_{1}z_{2})=g^{*}(X)$, thus $g^{*}(z_{1},z_{2}\mkern 1.5mu{|}\mkern 1.5muX)=0$. Since $g^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ and $f^{*}\mathord{\upharpoonright}Y_{1}XZ_{1}$ are the same polymatroids, the $n,m$-cone is part of the $n,1$-cone, as claimed. ∎
@@ 4.4. Problem Parameters
By Claim 6, the Maximum Entropy Method does not yield new inequalities when $f\mathord{\upharpoonright}X$ is modular. This is certainly the case when $|X|=1$, so we must have $|X|\geqslant 2$. By Claim 8, if $|Y|=|Z|=1$, then beyond the $1,1$-copy, no additional inequalities are generated. The smallest parameter setting when new entropy inequalities are expected as the number of copies grows is $|Y|=2$, $|X|=2$, and $|Z|=1$. We fix these sizes, as well as the labels of the members of each set as
@@
    Fix the number of copies $n$, called a generation. Determine the generating matrix $M$ of the cone $\mathcal{Q}^{*}$ as specified in Claim 2 using only polymatroids that satisfy the conditions of Claim 4, and are symmetric by Claim 5.
@@ 5. Computation
The cone $\mathcal{Q}^{*}$ whose non-Shannon bounding facets provide the new entropy inequalities sits in the $d_{1}=19$-dimensional Euclidean space with coordinates indexed by the non-empty subsets of $Y\mkern-2.0muX=\{abcd\}$ and $XZ=\{abz\}$. Fix the number of copies to $n\geqslant 1$. This choice also fixes the dimension of the vector $\mathbf{y}$ to $d_{3}=8\mkern 1.5mu{\cdot}\mkern 1.5mu4^{n}-1$. The generating matrix $M$ of the polyhedral cone $\mathcal{Q}^{*}$ from (2) is repeated here:
$\mathcal{Q}^{*}=\big\{\mathbf{x}\in\mathbb{R}^{d_{1}}:M\mkern 1.5mu{\cdot}\mkern 1.5mu(\mathbf{x},\mathbf{y})^{\top}\geqslant 0\mbox{ for some }\mathbf{y}\in\mathbb{R}^{d_{3}}\big\}.$ (4)
The modular part of $\mathcal{Q}^{*}$ is five-dimensional, and so its tight part sits in a $14$-dimensional subspace of $\mathbb{R}^{19}$, meaning that the dimension of $\mathbf{x}$ is reduced to $d_{1}=14$. By Claim 5, the polymatroid $f^{*}$ can be assumed to be symmetric for all permutations of the copies $Y_{i}=c_{i}d_{i}$. This means that the value $f^{*}(J)$ depends only on the cardinality of the sets
$\big\{i\leqslant n:J\cap Y_{i}=\{c_{i}\}\big\},~~~\big\{i\leqslant n:J\cap Y_{i}=\{d_{i}\}\big\},~~~\big\{i\leqslant n:J\cap Y_{i}=\{c_{i}d_{i}\}\big\},$
and on the set $J\cap XZ$. Since $|XZ|=3$, this property alone reduces $d_{3}$ to $8{n+3\choose 3}-1$. Additional conditional independences provided by Claim 4 reduce this number further to
$d_{3}=4{n\mkern 1.5mu{+}\mkern 1.5mu3\choose 3}+2{n\mkern 1.5mu{+}\mkern 1.5mu2\choose 2}+1,$ (5)
which number was verified empirically up to $n=10$.
A structural property of the polymatroid region $\Gamma_{\!abcd}$ on the four-element set abcdabcd allows us to further reduce the complexity of the polyhedral computation required in step (2) above. The region $\Gamma_{\!abcd}$ has a central part and six permutationally equivalent “protrusions,” depending on the signs of the Ingleton expressions
@@
If all of them are non-negative, then the restriction $f\mathord{\upharpoonright}abcd$ is a linear polymatroid; otherwise exactly one of these Ingleton expressions is negative, see, e.g., [25]. Accordingly, the cone $\mathcal{Q}^{*}$ is cut into seven parts by these Ingleton hyperplanes: the central part where all Ingleton values are non-negative, and six other parts where exactly one of the expressions is negative. The facets of each part of $\mathcal{Q}^{*}$ can be computed separately.
Parts of $\mathcal{Q}^{*}$ on the negative side of $\mkern 1.5mu], [adbc][adbc$], $bcad\mkern 1.5mu]$, and $[bdac]$ are isomorphic because swapping $a\leftrightarrow b$ and swapping $c\leftrightarrow d$ are symmetries of $\mathcal{Q}^{*}$. Therefore, it suffices to consider only one of them. The central part, where every Ingleton expression is non-negative, does not yield new inequalities. This follows from Lemma 9 below, as the elements of the central part are linear. Note that the proof uses the additional fact that $XZ$ has exactly three elements.
@@
Since every polymatroid on three elements is linear, and linearly representable polymatroids on three or four elements are representable over any field, we can assume, after scaling and using continuity, that both $f\mathord{\upharpoonright}abcd$ and $f\mathord{\upharpoonright}abz$ are $\mathbb{F}$-linearly representable over the same finite field $\mathbb{F}$. Denote the two representing vector spaces by $V^{1}$ and $V^{2}$, and consider the subspace arrangements $(V^{1}_{a},V^{1}_{b})$ and $(V^{2}_{a},V^{2}_{b})$ in the two vector spaces. Now $V^{i}_{a}$ and $V^{i}_{b}$ have dimensions $f(a)$ and $f(b)$, respectively, and their linear span has dimension $f(ab)$. Therefore, these arrangements are isomorphic, and $V^{1}$ and $V^{2}$ can be glued along the linear span of $(V^{1}_{a},V^{1}_{b})$ and $(V^{2}_{a},V^{2}_{b})$. This gluing yields an $\mathbb{F}$-linear polymatroid $g$ that has the same restrictions on abcdabcd and on abzabz as $f$ does. Since this $g$ is entropic, it has an $n$-copy for every $n\geqslant 1$. This $n$-copy is also an $n$-copy of $f$, as required. ∎
Consequently, up to the $a\leftrightarrow b$ and $c\leftrightarrow d$ symmetries, three mutually exclusive cases are left: $f[abcd\mkern 1.5mu]<0$, $f[acbd\mkern 1.5mu]<0$, and $f[cdab\mkern 1.5mu]<0$. Using the homogeneity of $\mathcal{Q}^{*}$, the Ingleton value can be set to 1-1, in effect taking a cross-section of $\mathcal{Q}^{*}$ that has one fewer dimension. Facets of the part of $\mathcal{Q}^{*}$ we are considering are also facets of these cross-sections; consequently, only facets of the cross-sections need to be computed. We consider these three cases separately in the subsections below.
@@
  • polymatroids have non-negative coordinates apart from the Ingleton coordinate.
The first property allows setting the Ingleton value explicitly. Based on the second property, the tight part of the cross-section can be separated by dropping some coordinates; and the third property potentially reduces the complexity of the polyhedral enumeration algorithm.